OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, 6) and (2, -6). Please show all of your work.

OpenStudy (anonymous):

first find the slope via \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{6+6}{-4-2}=\frac{12}{-6}=-2\]

OpenStudy (anonymous):

then use almighty point slope formula \[y-y_1=m(x-x_1)\] \[y-6=-2(x+4)\] \[y-6=-2x-8\] \[y=-2x-2\]